If sin(x)=(1/11) (in quadrant 1)& find sin(2x)=
1
Expert's answer
2011-08-03T11:31:18-0400
If sin x = 1/11,
cos x =(1 - sin2 x)1/2 = (1 - 1/121)1/2 =(120/121)1/2 = 2 √30 /11
sin 2x = 2 sinx cosx = 2*1/11*2√30 /11 = 4√30 / 121